# Testing stimulated emission photon directions

### v6

@@ Abstract
Keywords: lasers, photonics, STED microscopy, backward ASE, CPT symmetry, ring laser, synchrotron radiation, 2WQC
@@ I Introduction
Figure 1: Top: recent widely commented experiment by Steinberg group [1] has observed medium response (measured phase of 85Rb) before and after the applied impulse, proposing ”photon can spend a negative amount of time in an atom cloud” type of explanation, which does not seem physical.
Bottom: instead, as discussed here, it might be a result of just stimulated emission due to EM coupling of laser resonator and atoms as tiny resonators, required as some of the atoms were excited by magneto-optical trap (MOT) lasers, and if being CPT analog of absorption, should have opposite delay sign.
CPT symmetry (of charge conjugation (C) + parity transformation (P) + time reversal (T)) is shown [2] to be necessary for local Lorentz invariant quantum field theories, hence is generally believed to be satisfied by our nature. It has dozens of successful experimental confirmations [3], however, it seems all tests so far were microscopic - come from observations of single particles. Conscious macroscopic test was not found in literature, important especially due to T time symmetry which might lead to new applications - some are proposed and discussed here. For example through system response before impulse, like recently experimentally observed by Steinberg group in the news [1]111 www.scientificamerican.com/article/evidence-of-negative-time-found-….
@@
Figure 3: In commercially available e.g. erbium-doped optical fiber amplifiers (EDFA, or simpler semiconductor SOA, etc.) [4] there is observed strong backward ASE (amplified spontaneous emission) - increased emission from pumped doped fiber toward laser, exactly as in the sought CPT analog of ”laser causes excitation” effect. Top:Analogously to STED-like test, we can verify this photon direction by measuring delay between the impulse and reduced population of excited atoms of external target, e.g. by monitoring its spontaneous emission. For this purpose, the shown standard forward optical isolator should be removed, or better rotated to backward position. As erbium has very long millisecond-scale lifetimes, there is required fast photodiode. Fig. 7 contains further simplified test. Bottom: simplified universal setting as in [5], more practical to measure backward ASE delay, separation, transparency, and for their potential applications.
@@
As there are at least these two possibilities for stimulated emission photon directions, the final judge should be experiment. Related tests were earlier proposed using free electron laser [7] (earlier 2009 post), or ring laser [8]. However, they turned out too difficult to organize. This article simplifies test setting to STED-like, and later to optical amplifier - determining photon direction through delay. Hence it should be organized soon - to update this article, and depending on the outcome: leading to further discussed novel applications, or indicating macroscopic violation of CPT symmetry.
@@
Multiple updates of this article were mainly motivated by finding additional indirect experimental evidence in the literature (e.g. for Fig. 1, 3, 5), transparency with textbook view as consequent, convenient optical amplifier realization, and adding multiple potential applications. Also, QInter of our 2WQC group (https://www.qaif.org/2wqc): Vysakh Saji has recently performed the proposed experiment, obtaining unofficial result as $\Delta t\approx 0.0082$ ns for $l\approx d$ setting, this way confirming validity of CPT symmetry in macroscopic scale, to be updated with official results.
@@ II STED microscopy based tests
While spontaneous emission photons should have a random direction, the question of concern is direction of stimulated emission photons: ()($\rightarrow$)( → ) forward ano/or ()($\leftarrow$)( ← ) backward? Textbooks assume forward, discussion here based on CPT symmetry suggests backward as fundamental, and forward as its consequence (Fig. 4). Therefore, both can be true, the question to verify is existence of backward ()($\leftarrow$)( ← ).
@@ III-B Optical cavity and delocalized standing EM wave coupling
Being different types of EM field configurations, their interactions might also differ, for example with sophisticated devices like optical isolator - specialized to block localized optical photons travelling in one direction, but not necessarily delocalized standing EM waves for coupling - what should be tested, e.g. with setting as in Fig. 7. Finding material which acts as a mirror for optical photons, but is semi-transparent for delocalized EM waves for coupling, would be perfect to close resonator of potential backward lasers: only amplifying target deexcitation, e.g. for radiotherapy as in Fig. 10.
@@ III-C Rabi cycle as coupling, its asymmetrization
Bottom: proposed test if optical isolator blocks this EM coupling - monitoring backward ASE: amplified emission from erbium toward laser, for 3 optical isolator cases. It is observed without optical isolator (’none’ case), backward ASE photons are often blocked by $\rightarrow$ forward placed optical isolator. The main question is if $\leftarrow$ backward placed optical isolator leaves backward ASE - suggested both by CPT symmetry (switching photon and optical isolator direction), and coupling through delocalized EM wave.
@@ IV Some potential new applications
If confirmed, the above considerations could lead to many new applications, especially for the currently neglected backward beam - optimistically of standard e.g. optical fiber equipment (resonator coupling) or ring lasers, pessimistically of synchrotron radiation (forward/backward sources), like:
  • Forward beam, e.g. from synchrotron radiation, could allow for state preparation of quantum computers to nearly completely excited $N_{2}\approx N$ as in [15], backward beam for its CPT analog (postparation) for 2WQC [8] as in Fig. 9 in theory allowing for up to exponential speedup and better stability ([28, 29]).
@@
  • Backward beam might be transparent for $N_{2}=0$ medium, what could allow for much better penetration than forward (transparent for $N_{1}=0$) - e.g. for backward camera as in Fig. 11, various scanners, communication through obstacles. Finding material being mirror for optical photons, but semi-transparent for standing EM waves for coupling, it would be perfect to close resonator of potential backward laser.
  • Backward beam to stimulate specific chemical processes: releasing photons of chosen energy (and polarization), e.g. for chemical/technological processes, energy production, or medical therapies as in Fig. 10 - e.g. locally reduce lifetime of ATP molecule (stimulate decay also releasing this energy locally heating), or inhibit biochemical processes by amplifying deexcitation of crucial molecules, for example with multiple beams intersecting in cancer tissue, especially if better transparency is confirmed.
  • Backward camera maybe exploiting transparency and/or opposite delay e.g. for astronomical observations - focused on target’s backward beam, while standard camera is focused on its forward beam. It would replace CMOS/CCD matrix behind the lens, with matrix/surface of continuously excited approximately 3-level medium like dye, plus some monitoring of its population map - e.g. through spontaneous emission, or EIT (electromagnetically induced transparency).
@@
  • Backward beam for observations actively increasing deexcitation rate of excited objects, e.g. for thermal imaging.
@@
Figure 9: Original motivation: (example of) two-way quantum computer (2WQC) [8] as adding a process which in CPT perspective becomes the original state preparation process. For photonic QC state preparation is usually performed by laser impulse, suggesting to just add ’reversed’ impulse: delayed by propagation time, and behind backward placed optical isolator - this way from CPT perspective photon trajectories should be reversed. In comparison to standard 1WQC, in theory 2WQC allows for up to exponential speedup e.g. to solve NP problems, also better stability and error correction [28, 29]. Such classical 2WCC photonic computer on continuous beam might be also interesting.
Figure 10: Being able to split absorption and stimulated emission equations, for example with CPT analog of synchrotron radiation: beam resulting from reversed electron trajectories as in Fig. 6, or maybe standard laser with backward placed optical isolator as in Fig. 7, one of many new applications could be radiotherapy trying to locally amplify emission of photons of chosen energy and polarization. For example to try to locally increase degradation of energy carrying molecules like ATP (if having emission spectrum) - to starve cancer tissue, and heat up by releasing this energy. Intersecting multiple such beams, maybe additionally rotating them for fixed intersection point, we could make the effect there orders of magnitude stronger then for intermediate tissues, to make it negligible outside of targeted tissue. Different issues could be also targeted this way, e.g. trying to disable toxic molecules, inhibit chosen biochemical pathways by stimulating deexctiation of crucial molecules e.g. of pathogens, or just for local cooling, etc.
Figure 11: Top: Backward camera observing backward ASE caused by the target, especially interesting if confirmed better transparency and/or opposite delay sign. Can be e.g. based on (multi-mode) fiber with circulators as in [5] and Fig. 3, or more practical with semitransparent mirror and CCD/CMOS - both need some continuously pumped medium (maybe from some display technology). Alternatively one could monitor spontaneous emission of the pumped medium.
Bottom: Backward lighted camera: applying backward ASE effect on the target to amplify its emission in chosen spectrum toward camera (e.g. thermal also cooling), or on laser medium used to send quantum cryptography protocols like BB84, to increase its number of produced photons and steal/read some.

### v8

@@ Abstract
Keywords: lasers, photonics, STED microscopy, backward ASE, CPT symmetry, ring laser, synchrotron radiation, 2WQC, optical fiber, EDFA amplifier, optical isolator
@@ I Introduction
Figure 1: Top: recent widely commented experiment by Steinberg group [1] has observed medium response (measured phase of 85Rb) before and after the applied impulse, proposing ”photon can spend a negative amount of time in an atom cloud” explanation, which does not seem physical, however, theoretical explanation for this article uses forward/backward-in-time propagators [2].
Bottom: as discussed here, intuitive explanation of such forward/backward propagators, like in two-state vector formalism [3], could come from absorption/stimulated emission being CPT analogs, hence should have opposite delay.
CPT symmetry (of charge conjugation (C) + parity transformation (P) + time reversal (T)) is shown [4] to be necessary for local Lorentz invariant quantum field theories, hence is generally believed to be satisfied by our nature. It has dozens of successful experimental confirmations [5], however, it seems all tests so far were microscopic - come from observations of single particles. Conscious macroscopic test was not found in literature, important especially due to T time symmetry which might lead to new applications - some are proposed and discussed here. For example through system response before impulse, like recently experimentally observed by Steinberg group in the news presented in Fig. 1 [1]111 www.scientificamerican.com/article/evidence-of-negative-time-found-….
@@
Figure 3: In commercially available e.g. erbium-doped optical fiber amplifiers (EDFA, or simpler semiconductor SOA, etc.) [6] there is observed strong backward ASE (amplified spontaneous emission) - increased emission from pumped doped fiber toward laser [7], exactly as in the sought CPT analog of ”laser causes excitation” effect, usually removed by forward optical isolators. Top: Analogously to STED-like test, we can verify this photon direction by measuring delay between the impulse and reduced population of excited atoms of external target, e.g. by monitoring its spontaneous emission. For this purpose, the shown standard forward optical isolator should be removed, or better rotated to backward position. Bottom: simplified setting from [7] seems much more practical for tests: of delay (opposite if CPT analog), separation (leaving forward photons in CPT perspective), transparency (blocked by excited/deexcited?) and potential future applications.
@@
As there are at least these two possibilities for stimulated emission photon directions, the final judge should be experiment. Related tests were earlier proposed using free electron laser [9] (earlier 2009 post), or ring laser [10]. However, they turned out too difficult to organize. This article simplifies test setting to STED-like, and later to fiber optics amplifiers - determining photon direction through delay. Hence it should be organized soon - to update this article, and depending on the outcome: leading to further discussed novel applications, or indicating macroscopic violation of CPT symmetry.
@@
Multiple updates of this article were mainly motivated by finding additional indirect experimental evidence in the literature (e.g. for Fig. 1, 3, 5), transparency with textbook view as consequent, convenient optical amplifier realization, and adding multiple potential applications. Also, QInter of our 2WQC group (https://www.qaif.org/2wqc): Vysakh Saji has performed the proposed experiment, obtaining looking confirming results, but without followup. Recently we have started local tests (with Tomasz Urbańczyk and Tomasz Kawalec), which lead to observation/difficulty that nearly all available laboratory lasers have built-in optical isolator, what seems one of reasons backward ASE properties are practically unknown, also confirmed by discussion with multiple photonics and fiber optics experts. This article should be updated with further tests.
@@ II STED microscopy based tests
While spontaneous emission photons should have a random direction, the question of concern is direction of stimulated emission photons: ()($\rightarrow$)( → ) forward ano/or ()($\leftarrow$)( ← ) backward? Textbooks assume forward, discussion here based on CPT symmetry suggests backward as fundamental, and forward as its consequence (Fig. 4). Therefore, both can be true, the question to verify is existence of backward ()($\leftarrow$)( ← ). All tests require laser without built-in forward optical isolator.
@@ III-B Optical cavity and delocalized standing EM wave coupling
Being different types of EM field configurations, their interactions might also differ, for example with sophisticated devices like optical isolator - specialized to block localized optical photons travelling in one direction, but not necessarily delocalized standing EM waves for coupling - what should be tested, e.g. with setting as in Fig. 7. As in Fig. 9, some optical isolators could be enhanced with mirrors to be reflective instead of blocking in the opposite direction, useful e.g. to enclose resonator of potential backward lasers: only amplifying target deexcitation, e.g. for radiotherapy as in Fig. 12.
@@ III-C Rabi cycle as coupling, its asymmetrization
Bottom: turns out nearly all lasers have built-in optical isolator, what might prevent required Rabi bidirectional energy exchange like for FEL. Searching literature, while for microwaves Rabi cycle has straightforward realization, for lasers there are used tricks like added resonator [29], or Mach-Zehnder type setting ([30, 31]), requiring analysis like in ”Asking photons where they have been” [32]. Just removing this isolator might allow for simpler realizations.
@@ IV Some potential new applications
The above considerations could lead to many new applications, especially if confirmed the suggested by CPT symmetry properties of discussed backward beam like in backward ASE [7]: possibility to separate from forward, better transparency, and opposite delay. Optimistically it can be obtained by laser with backward directed optical isolator, or maybe ring laser. Pessimistically synchrotron radiation would be needed (forward/backward sources) as in Fig. 6. For example:
  • Forward beam, e.g. from synchrotron radiation, could allow for state preparation of quantum computers to nearly completely excited $N_{2}\approx N$ as in [17], backward beam for its CPT analog (postparation) for 2WQC [10] as in Fig. 11, in theory allowing for up to exponential speedup and better error correction ([34, 35]).
@@
  • Improvements of current technologies, like STED microscope not photobleaching the sample: by reversing optical isolator of depletion laser, or direct laser realization of Rabi cycle by removing optical isolator (Fig. 7).
  • Backward beam in theory should be transparent for $N_{2}=0$ medium, what could allow for much better penetration than forward (transparent for $N_{1}=0$) - e.g. for radiotherapy as in Fig. 12, backward camera as in Fig. 10 and various scanners e.g. CAT for 3D map of $N_{2}$ atoms/molecules as in Fig. 13, communication through obstacles, etc.
  • Backward beam to stimulate specific chemical processes: releasing emission spectrum photons of chosen energy (and polarization), e.g. for chemical/technological processes, energy production, or medical therapies as in Fig. 12 - e.g. locally reduce lifetime of NADH molecule to starve cancer tissue and release this energy as thermal, or inhibit biochemical processes by amplifying deexcitation of crucial molecules or injected contrast medium, for example with multiple rotating beams intersecting in cancer tissue, especially if better transparency is confirmed.
  • Backward camera as in Fig. 10 maybe exploiting transparency and/or opposite delay e.g. for astronomical observations (for example pulsars could produce forward+backward synchrotron radiation) - focused on target’s backward beam (or transmission of added backward beam), while standard camera is focused on its forward beam. It would replace CMOS/CCD matrix behind the lens, with matrix/surface of continuously excited approximately 3-level medium like dye or erbium, plus some monitoring of its population map - e.g. through spontaneous emission, or EIT (electromagnetically induced transparency). Pairing with backward beam in line, could allow for various scanning (e.g. medical, material science, geological) allowing (emission) CAT scanning for 3D maps of $N_{2}$ excited atoms/molecules between them as in Fig. 10, 13.
    Figure 9: While most laboratory lasers have built-in (forward) optical isolator - can be referred as forward laser (only pushing photons), especially for Rabi cycle as in Fig. 7 it is worth to distinguish (bidirectional or) just laser having removed optical isolator. For various mentioned applications it is also worth to consider backward laser as CPT transformed forward laser: with reversed optical isolator, should only pull photons. Polarization independent optical isolators usually change angle for the opposite photon direction, allowing to add mirrors inside to reflect back such photons (e.g. extending laser resonator toward one time direction) - mirrored backward laser with isolator containing such mirrors should be much more energy efficient, might also improve forward laser as having lower energy loss toward $-t$. Angle, polarization, path for to reflect such reverse photons would need optimization, similar effect could be realized with circulator, ideally would be getting such effect with metamaterials.
    Figure 10: Top: Backward camera/detector observing backward ASE caused by the target (or added backward laser), especially interesting if confirmed better transparency and/or opposite delay sign. Detector can be e.g. based on (multi-mode) fiber with circulators as in [7] and Fig. 3. For camera there can be used semitransparent mirror and CCD/CMOS - both need some continuously pumped medium (maybe from some display technology). Alternatively one could monitor spontaneous emission of the pumped medium, e.g. with camera behind. Pairing backward detector/camera with pointing backward laser could allow for transmission scanning for up to 3D maps of excited molecules between (e.g. medical, material science, geological scanning) - like in standard transmission scanning, but of emission spectrum instead of absorption.
    Bottom: Backward lighted camera: applying backward ASE effect on the target to amplify its emission in chosen spectrum toward camera (e.g. thermal also cooling), or on laser medium used to send quantum cryptography protocols like BB84, to increase its number of produced photons and steal/read some - e.g. going through control of CNOT gate to work on single photons.
@@
  • Backward-lighted camera as in Fig. 10: with added backward beam for observations actively increasing deexcitation rate of excited objects, e.g. for thermal imaging.
    Figure 11: Original motivation: (example of) two-way quantum computer (2WQC) [10] (and its classical hydrodynamical analog) as adding a process which in CPT perspective becomes the original state preparation process. For photonic QC state preparation is usually performed by laser impulse, suggesting to just add ’reversed’ impulse: delayed by propagation time, and behind backward placed optical isolator - this way from CPT perspective photon trajectories should be reversed, like improving two-way flow control in hydrodynamical analog. In comparison to standard 1WQC, in theory 2WQC allows for up to exponential speedup e.g. to solve NP problems, also better stability and error correction [34, 35]. Such classical 2WCC photonic computer on continuous beam or impulses seems also worth considering.
    Figure 12: If being able to separate absorption and stimulated emission equations, for example with CPT analog of synchrotron radiation: beam resulting from reversed electron trajectories as in Fig. 6, or maybe standard laser with backward placed optical isolator as in Fig. 9, one of many new applications could be radiotherapy trying to locally amplify emission of photons of chosen energy and polarization. For example to try to locally increase degradation of energy carrying molecules like NADH (nicotinamide adenine dinucleotide) crucial for cell energy metabolism, or other autofluorescent molecules with emission spectrum (lists e.g. in [37, 38, 39]) - to starve cancer tissue, and heat up by releasing this energy. Intersecting multiple such beams, or e.g. one beam spread in 1D by two cylindrical lenses, additionally rotating them for fixed focus intersection point, we could make the (usually nonlinear) effect there orders of magnitude stronger then for intermediate tissues, to make it negligible outside of targeted tissue. Different issues could be also approached this way, e.g. contrast agent as in photodynamic therapy [40], or trying to disable toxic molecules (like $\sim$1280nm for singlet oxygen [41]), inhibit chosen biochemical pathways by stimulating deexctiation of crucial molecules e.g. of pathogens, etc.
@@
Figure 13: Simplified Fig. 10 diagram for measurement of (integrated over light path) absorption coefficient, which can be extended e.g. to CAT scanner creating 3D maps of absorption coefficient from 2D projections in multiple angles, reconstructed through Radon transform [42]. Above CPT symmetric setting suggests how to analogously measure emission coefficients, what could be extended to its CAT scanning. For medical applications e.g. of autoluminescent molecules (gaining energy from cell chemistry) to make 3D maps of e.g. tryptophan ($\sim$340nm), collagen ($\sim$360nm), elastin ($\sim$410nm), NADH ($\sim$460nm), flavins ($\sim$525nm) [37], or some contrast agent - especially if possible to separate, and indeed being blocked only by molecules emitting in such chosen frequency (better transparency). Using optimized multiple frequencies, could allow for noninvasive detailed (high spatial and temporal resolution) 3D maps of tissues with detected pathologies like cancer. The source could be laser with reversed optical isolator, or ring laser, or synchrotron radiation. The detector could be backward camera as in Fig. 10: pumped surface (e.g. screen) with monitored population, or adding e.g. half-silvered mirror and standard detector.